If 300cm3 of hydrogen diffuses through a porous pot in 50seconds. How long will it take 500cm3 of hydrogen to diffuse through the same pot. (Hydrogen=1,oxygen=16)



1
Expert's answer
2021-06-03T09:16:27-0400
"\frac{300t}{500(50)}=\sqrt{\frac{1}{16}}\\t=20.8\ s"